# Variability in the massive black hole binary candidate SDSS J2320+0024: no evidence for periodic modulation

### v1

@@ Appendix A Analysis on the r-band
Figure 4: Corner plot of the posteriors of the different models. DRW is blue, modified DRW is red, QPO is green, and Periodic is orange. Contours are drawn at a level to include 90% of the posterior probability, while the black dashed lines refer to the period reported in Fatović et al. (2023). For reasons of clarity, although the fit is performed on standardized data, we report the marginal posterior distribution of the period in units of $\log_{10}$ Days. Please note the different range of $\log_{10}{T}$ in the corner plot. No short ($\simeq 260$ days) oscillations are identified by the nested sampling algorithm.
Appendix B Quality of the fit
@@
The analysis presented in this paper is based on the Bayes ratio that, as already discussed in the paper, is the most robust approach to model selection and model comparison. However, this statement is true only under the assumption that all the models being compared are a reasonably good description of the data. To check this, we report the standardized residuals (i.e., (model-data)/error) for the same models presented in Fig. 1, testing whether they are consistent with a normal distribution of zero mean and unit dispersion. For each model, we provide the residuals using either the median model (blue) or the model maximizing the likelihood (red), also reporting the mean ($\mu$) and the standard deviation ($\sigma$) of the residuals.
@@
In the case of good modelling of the data, the standardized residuals are expected to follow the normal distribution with $\mu=0$ and $\sigma=1$. To quantitatively asses this we performed a Kolmogorov-Smirnov test (Smirnov 1948) founding that the DRW and the Modified DRW residuals always follow the expected distribution (p-value larger than the commonly adopted threshold of 0.05), the Periodic model residuals are never normally distributed while the hypothesis of a normal distribution for the QPO model can be accepted only in the case of the maximum likelihood model being compared with the data. Notably, even in this case, the p-value of the QPO model is smaller than that of the DRW or Modified DRW model.
@@
Figure 5: Histograms of the residuals (model-data)/error for the models shown in Fig. 1. Each plot also reports the mean, standard deviation, and the p-values of a Kolmogorov–Smirnov test. The color coding follows the one reported in Fig. 1: blue for the residual of the median model, while red for the model which maximizes the likelihood. From top to bottom, we show the DRW, the modified DRW, the QPO, and the Periodic models.

### v2

@@ Appendix A Analysis on the r-band
Figure 4: Corner plot of the posteriors of the different models. DRW is blue, modified DRW is red, QPO is green, and Periodic is orange. Contours are drawn at a level to include 90% of the posterior probability, while the black dashed lines refer to the period reported in Fatović et al. (2023). For reasons of clarity, although the fit is performed on standardised data, we report the marginal posterior distribution of the period in units of $\log_{10}$ Days. Please note the different range of $\log_{10}{T}$ in the corner plot. No short ($\simeq 273$ days) oscillations are identified by the nested sampling algorithm.
Appendix B Analysis on the i-band
Here we repeat the same analysis performed in Sec. 3 and Appendix A focusing on i-band data. The best-fit light curves for the four models are plotted with their uncertainties in Fig. 5, while the parameters describing the models are reported in Tab. 3.  [^8] In this case, we rescale the data using a mean i-band magnitude of 21.131 (standard deviation of 0.318) and a mean MJD of 54963.152 (standard deviation of 2588.802).
As for the g- and r-bands, the analysis prefers the DRW, Modified DRW, and QPO models to the periodic one, with a slight preference for the DRW. The posterior distribution of the period found in the Periodic model becomes bimodal (see Fig. 6), explaining the large uncertainties in the predictive posterior (see Fig. 5 panel d), with the smallest (and highest likelihood) period ($\simeq 448$ days), being closer, but still larger, to that reported by Fatović et al. (2025) than the corresponding highest peaks found for the g- and r-bands.
As commented in Appendix A, the lower evidence for the periodic model and, within the periodic model, the lack of a consistent period in adjacent bands and the bimodal distribution of the periods further strengthen our conclusion about the lack of periodicity in the current data of J2320+0024.
Table 3: Summary of the different model parameters for the fit to the i-band light curve of SDSS J2320+0024.  [^9] From left to right, the columns provide a brief description of each parameter, its reference name as used in this work, the assumed prior range, and the best-fit values with their credibility intervals for all the different models. The first table block reports the parameters estimated as the median, 16th, and 84th percentiles of the posterior distribution, while the second table block reports the set of parameters that maximise the likelihood. All the priors are log-uniform, and all the parameters are in dimensionless units since the data have been standardised before fitting. The last row reports the Bayesian log evidence of each model clearly disfavoring the Periodic model.
Description Name Prior range DRW Modified DRW QPO Periodic
50th, 16th, 84th Percentiles
Scale length $\log_{10}{\tau}$ [3,1.7][-3,1.7] $1.2^{+0.3}_{-0.3}$ $1.2^{+0.3}_{-0.4}$ $1.2^{+0.4}_{-0.3} -$
Dispersion $\log_{10}{\sigma}$ [3,1.7][-3,1.7] $0.6^{+0.2}_{-0.1}$ $0.6^{+0.2}_{-0.2}$ $0.6^{+0.2}_{-0.1}$ $0.1^{+0.3}_{-0.2}$
Slope $\gamma$ $[0.5,5]$ - $1.2^{+0.7}_{-0.5}$ - -
Period $\log_{10}{T}$ [2,2][-2,2] - - $0.6^{+1}_{-0.9}$ $0.15^{+0.01}_{-0.6}$
Maximum Likelihood
Scale length $\log_{10}{\tau}$ [3,1.7][-3,1.7] 1.19-1.19 1.16-1.16 1.12-1.12 -
Dispersion $\log_{10}{\sigma}$ [3,1.7][-3,1.7] $0.63$ $0.90$ $0.58$ 0.16-0.16
Slope $\gamma$ $[0.5,5]$ - $2.1$ - -
Period $\log_{10}{T}$ [2,2][-2,2] - - 0.21-0.21 0.76-0.76
log evidence $\log{Z}$ - -73.0 -74.1 -73.5 -124.3
Figure 5: Best-fit for the i-band light curve of J2320+0024. The black points represent the data with their errors, the blue solid line and the shaded area represent the median, 16th, and 84th percentiles from the posterior distributions, and the red dashed line refers to the maximum likelihood model. From top to bottom: DRW (a), modified DRW (b), QPO (c), and Periodic (d), we refer to Tab. 1 for the estimated best-fit parameters.
Figure 6: Corner plot of the posteriors of the different models. DRW is blue, modified DRW is red, QPO is green, and Periodic is orange. Contours are drawn at a level to include 90% of the posterior probability, while the black dashed lines refer to the period reported in Fatović et al. (2023). For reasons of clarity, although the fit is performed on standardised data, we report the marginal posterior distribution of the period in units of $\log_{10}$ Days. Please note the different range of $\log_{10}{T}$ in the corner plot. No short ($\simeq 273$ days) oscillations are identified by the nested sampling algorithm.
Appendix C Quality of the fit
@@
The analysis presented in this paper is based on the Bayes ratio that, as already discussed in the paper, is the most robust approach to model selection and model comparison. However, this statement is true only under the assumption that all the models being compared are a reasonably good description of the data. To check this, we report the standardised residuals (i.e. (model-data)/error) for the same models presented in Fig. 1, testing whether they are consistent with a normal distribution of zero mean and unit dispersion. For each model, we provide the residuals using either the median model (blue) or the model maximizing the likelihood (red), also reporting the mean ($\mu$) and the standard deviation ($\sigma$) of the residuals.
@@
In the case of good modelling of the data, the standardised residuals are expected to follow the normal distribution with $\mu=0$ and $\sigma=1$. To quantitatively asses this we performed a Kolmogorov-Smirnov test (Smirnov 1948) founding that the DRW and the Modified DRW residuals always follow the expected distribution (p-value larger than the commonly adopted threshold of 0.05), the Periodic model residuals are never normally distributed while the hypothesis of a normal distribution for the QPO model can be accepted only in the case of the maximum likelihood model being compared with the data. Notably, even in this case, the p-value of the QPO model is smaller than that of the DRW or Modified DRW model.
@@
Figure 7: Histograms of the residuals (model-data)/error for the models shown in Fig. 1. Each plot also reports the mean, standard deviation, and the p-values of a Kolmogorov–Smirnov test. The color coding follows the one reported in Fig. 1: blue for the residual of the median model, while red for the model which maximises the likelihood. From top to bottom, we show the DRW, the modified DRW, the QPO, and the Periodic models.